# Semiring Activation in Neural Networks

### v1

@@ 4.3 Initialization
$\operatorname*{Var}(\overline{x}_{j})\approx\operatorname*{Var}\left(\sum_{i=1}^{n}\frac{\partial y_{i}}{\partial x_{j}}(0,\ldots,0)\overline{y}_{i}\right)=n\operatorname*{Var}(\overline{y})\sum_{i=1}^{n}\left(\frac{e^{\mu w_{ij}}}{\sum_{k=1}^{n}e^{\mu w_{ik}}}\right)^{2},$
@@
$\sum_{i=1}^{n}\left(\frac{e^{\mu w_{ij}}}{\sum_{k=1}^{n}e^{\mu w_{ik}}}\right)^{2}\approx\frac{1}{n}.$ (10)
This condition is exactly satisfied if all the weights are constant, say $w_{ij}=K\in\mathbb{R}$ for all $i=1,\ldots,n$ and $j=1,\ldots,m$, then
$\sum_{i=1}^{n}\left(\frac{e^{\mu w_{ij}}}{\sum_{k=1}^{n}e^{\mu w_{ik}}}\right)^{2}=n\left(\frac{e^{\mu K}}{n\,e^{\mu K}}\right)^{2}=\frac{1}{n}$
Satisfying both the forward condition (9) and backward condition (10) is not possible. However, the fair tropical initialization scheme does present a reasonable trade-off between these two clashing requirements. Indeed if we initialize using this scheme (assuming $n>m$ for the moment) we get a matrix like (6). where we let $K>0$ if $\mu>0$ and $K<0$ if $\mu<0$. On the one hand, most entries have the same $-K$ value and so satisfy (10) to some degree. On the other hand, in every row we have a single element (the zero values) that are relatively dominant with respect to the other elements in the row, thus catering to (9). Hence, we will use the tropical fair initialization scheme for the logarithmic operators as well.

### v2

@@ 4.3 Initialization
$\operatorname*{Var}(\overline{x}_{j})\approx\operatorname*{Var}\left(\sum_{i=1}^{n}\frac{\partial y_{i}}{\partial x_{j}}(0,\ldots,0)\overline{y}_{i}\right)=\operatorname*{Var}(\overline{y})\sum_{i=1}^{n}\left(\frac{e^{\mu w_{ij}}}{\sum_{k=1}^{n}e^{\mu w_{ik}}}\right)^{2},$
@@
$\sum_{i=1}^{n}\left(\frac{e^{\mu w_{ij}}}{\sum_{k=1}^{n}e^{\mu w_{ik}}}\right)^{2}\approx 1.$
or
$\sum_{i=1}^{n}\left(e^{\mu w_{ij}}\right)^{2}\approx\left(\sum_{i=1}^{n}e^{\mu w_{ij}}\right)^{2}$ (10)
for all $j=1,\ldots,m$. We can satisfy this condition exactly by choosing a single $I_{j}=1,\ldots,n$ for each $j=1,\ldots,m$ and set that weight to $w_{I_{j},j}=0$ and set the other weights to $w_{ij}=-\operatorname*{sgn}(\mu)\infty$. In that case $e^{\mu w_{ij}}=1$ if $i=I_{j}$ and $e^{\mu w_{ij}}=0$ if not and consequently both sides of (10) are equal to $1$.
[…]
Consider the loss gradient for an individual parameter:
$\overline{w}_{ij}=\frac{e^{\mu(x_{j}+w_{ij})}}{\sum_{k=1}^{m}e^{\mu(x_{k}+w_{ik})}}\overline{y}_{i}.$ (11)
If we try to satisfy (9) and (10) then we would set $w_{ik}=0$ for some $k=1,...,n$ for every $i=1,\ldots,n$. This implies that the denominator in (11) is greater than or equal to $1$. If we subsequently set $w_{ij}=-\operatorname*{sgn}{(\mu)}K$ for $j\neq k$ and $K$ a very large number then the numerator of (11) becomes vanishingly small, possible zero in floating point format. This essentially freezes the parameter’s value since any updates applied to it would become practically zero.
Assuming the inputs $(x_{j})_{j=1}^{m}$ are centered, the expected value of the fraction in (11) is given by
$\frac{e^{\mu w_{ij}}}{\sum_{k=1}^{m}e^{\mu w_{ik}}},$ (12)
which gives values in the range $[0,1]$ for any $i=1,\ldots,n$ and $j=1,\ldots,m$. We can avoid this fraction becoming (effectively) zero from the start by initializing all the parameters to roughly the same value, i.e. for all $i=1,\ldots,n$ set
$w_{i1}\approx w_{i2}\approx\ldots\approx w_{im}.$ (13)
In this case we achieve the maximum of the minimum where (12) gives the value $\nicefrac{{1}}{{m}}$ for all $i=1,\ldots,n$ and $j=1,\ldots,m$.
Satisfying the forward condition (9), backward condition (10) and parameter trainability condition (13) all at the same time is not possible. However, the fair tropical initialization scheme does present a reasonable trade-off between these three clashing requirements. Indeed if we initialize using this scheme (assuming $n>m$ for the moment) we get a matrix like (6), where we let $K>0$ if $\mu>0$ and $K<0$ if $\mu<0$. First, most entries have the same $-K$ value and so satisfy (13) to some degree. Second, in every row we have a single element (the zero values) that are relatively dominant with respect to the other elements in the row, thus catering to (9). Third, in every column we have at most $\lceil\nicefrac{{n}}{{m}}\rceil$ zero values that are relatively dominant with respect to the other elements in the column, thus catering to (10).